calculate the work done by 0.1 mole of a gas at 27°C to double its volume at constant pressure R = 2 cal mol^-1 K^-1

Grade:11

1 Answers

Arun
6 years ago
Dear student
W = p * \Delta V
W = p * (2V - V) = p* V
Now
W = - p* V = - n * R * T
= 0.1 * 2 * 300 = 60 cal
Regards
Arun (askIITians forum expert)